Please help me to solve y^3-4y^2-64y+256=0
Answer by tutorcecilia(2152)

y^3-4y^2-64y+256=0 [factor by grouping]
(y^3-4y^2)-(64y+256)=0
y^2(y-4)-64(y-4)
(y^2-64)(y-4)=o [set each factor equal to zero and solve]
y^2-64=0
sqrt(y^2)=sqrt(64)
y=+-8
and y-4=0
y=4
.
check by plugging each value of y back into the formula and solve. the answer should be zero.